OpenStudy (anonymous):

guys, as k->infinity, doesn't (1+1/k)^k tend toward 1, not e?

OpenStudy (anonymous):

Whoops, nevermind... I was evaluating the limit of 1+1/k only, then raising it to the k power. My bad

OpenStudy (phi):

In case you are interested in this stuff, here is Khan's videos on compounding and e https://www.khanacademy.org/math/algebra2/exponential_and_logarithmic_func/continuous_compounding/v/introduction-to-compound-interest-and-e

OpenStudy (amoodarya):

\[(1+\frac{1}{k})^k=1+\left(\begin{matrix}k \\ 1\end{matrix}\right)1^{k-1}(\frac{1}{k})^1+\left(\begin{matrix}k \\ 2\end{matrix}\right)1^{k-2}(\frac{1}{k})^2+...=\\1+\frac{k}{k}+\frac{k(k-1)}{2k^2}+...=\\2+\frac{(k-1)}{2k}+...\] obviously we have \[2<2+\frac{(k-1)}{2k}+...\]

OpenStudy (amoodarya):

if you try to find upper bound you will find \[2+\frac{(k-1)}{2k}+...<3\]

OpenStudy (amoodarya):

so \[2< (1+\frac{1}{k})^k<\]

OpenStudy (amoodarya):

in fact \[\lim_{k \rightarrow \infty} (1+\frac{ 1 }{ k })^k \rightarrow e\]